OpenStudy (anonymous):

(08.03)Solve the system of equations and choose the correct answer from the list of options. x − y = 7 y = 3x + 12 2 over 19 comma 2 over 33 negative 2 over 19 comma negative 33 over 2 negative 19 over 2 comma negative 33 over 2 19 over 2 comma 33 over 2

OpenStudy (anonymous):

someone help please

OpenStudy (haseeb96):

There are two equations We call first equation as eq#1 and 2nd equation as eq#2 x-y=7.................eq#1 y=3x+12 .............eq#2 now what you have to do Just plug y's value of eq#2 into equation 1 at the place of y you will get this one x-3x-12 =7 -2x = 19 x= -19/2 now put x= -19/2 into eq#1 and get the value of y -19/2 -y=7 y=-19/2 -7 y=-19-14/2 y=-33/2 so your (x,y)= (-19/2, -33/2)

OpenStudy (haseeb96):

so, C is correct answer

OpenStudy (anonymous):

Haseeb96 is correct